You want to obtain a sample to estimate a population mean. Based on previous evidence, you believe the population standard deviation is approximately σ=61.9σ=61.9. You would like to be 90% confident that your estimate is within 1.5 of the true population mean. How large of a sample size is required?
Answer #1

population standard deviation, σ=61.9

Critical value , Z_{0.1/2} = 1.6449

Let n be the required sample size

Margin of error, E = 1.5

Z_{0.1/2}\frac{\sigma }{\sqrt{n}} =1.5

1.6449 \frac{61.9}{\sqrt{n}} = 1.5

\sqrt{n} = \frac{1.6449 * 61.9}{1.5}

\sqrt{n} = 67.8

Squaring both sides

n = 4607.63 = 4608

So the required sample size is 4608.